In Newton's ring experiment the radius of curvature of the 'lens is 0.7 m and the diameter of the 10th bright ring formed due to reflected light is 0.5 cm. Calculate the wavelength of light used.
1
Expert's answer
2019-05-27T11:10:23-0400

Radii of the 10th bright ring:


"r^2_{10} = \frac{2\cdot 10 +1}{2} \lambda R \Rightarrow \lambda = \frac{2 r^2_{10}}{21R} = \frac{d^2_{10}}{42R}"

Where R = 0.7m.


"\lambda = \frac{5^2 \,\text{mm}^2}{42\cdot 0.7 \,\text{m}} = 0.85\cdot 10^{-6}\,\text{m} = 850 \,\text{nm}"